OpenStudy (anonymous):

How can you solve square roots by hand, such as \sqrt{5}? I'm looking for a method that doesn't require approximation and then checking back.

OpenStudy (anonymous):

you can get it by a recursive method but since this number is irrational you cannot get it using elementary operations. do you want a recursion?

OpenStudy (anonymous):

quickest method is mechanic's rule aka newton's method. make a guess, say 2, then take \[\frac{1}{2}(2+\frac{5}{2})\] rinse and repeat

OpenStudy (anonymous):

i.e. \[x_{n+1}=\frac{1}{2}(x_n+\frac{5}{x_n})\]

OpenStudy (anonymous):

i did see a "long division" style method once but it is annoying and painfully slow.

OpenStudy (anonymous):

I'm not looking for a guess and check...I feel those are better left for calculators. What is this long division style method you were talking about?